Let P be the point on the parabola y2=3x such that OP make an angle of pi/6 with the x-axis, where O is the origin. A normal is drawn to the parabola at P intersecting the axis of the parabola at Q. If S is the focus of the parabola, then S is equal to:

Grade:11

1 Answers

Aditya Gupta
3 years ago
hello meghana, note that there is some mistake in the question, as in the last line u have written:
If S is the focus of the parabola, then S is equal to
obviously we can easily calculate S by writing y2=3x as y2= 4*(3/4)x so that S= (3/4, 0).
kindly repost the correct ques with a pic attached